OpenStudy (surana):

What is the equation of the line which includes points (2,3) and (0,10)?

OpenStudy (hockeychick23):

ok so first step in the equation is to use point-slope form. Also known as (y2-y1)/(x2-x1). So its: 10-3/0-2= -7/2 next you choose one of the two sets of points to plug in as (x,y) 10=-7/2(0)+b 10= 0+b b= 10 now we have your y intercept. So your equation is y=-7/2x+10

OpenStudy (wolf1728):

(2,3) and (0,10) The equation would be of the form y = mx + b First we must calculate the slope. which can be determined by: m= 10 -3 / 0 -2 m = 7/-2

OpenStudy (surana):

Okay.

OpenStudy (surana):

So after calculating the slope, what's next?

OpenStudy (hockeychick23):

next you choose one of the two sets of points to plug in as (x,y) 10=-7/2(0)+b 10= 0+b b= 10 now we have your y intercept.

OpenStudy (hockeychick23):

i used (0,10)

OpenStudy (surana):

So after the Y intercept is found, I have to calculate the other one?

OpenStudy (hockeychick23):

nope. after you find your y-intercept your done. your equation is: y= -7/2x+10

OpenStudy (surana):

Oh. Thank you.

OpenStudy (hockeychick23):

anytime :)

OpenStudy (surana):

Thank you both for your help.

OpenStudy (wolf1728):

u r welcome (even though hockeychick actually got the answer)

OpenStudy (surana):

:D